1.1 calculate the break-even quantity. (3 marks) 1.2 calculate the target sales value using the marginal income ratio, if a profit of r600 000 is desired. (3 marks) 1.3 the sales manager made the following proposal to increase profitability: decrease the selling price by r40 per unit and increase advertising expense by r24 000 with the expectation that sales volume will increase by 10%. should the sales manager’s proposal be accepted? motivate your answer with the relevant calculations.
1
Expert's answer
2020-04-22T16:58:53-0400

1.1Break-even point can be determined by the following methods in physical terms and in monetary terms.

natural

Fixed costs + Marginal revenue per unit

where the marginal income per unit is the difference between price and variable costs per unit of product

monetary value

Fixed expenses + Margin income ratio

where the marginal income ratio is the ratio of marginal income per unit to the price


FC=282 000+150 000=432 000

"VC=270+180+90+0.1\times(3000\times900)=540+270 000=270540"

"VC unit=\frac{270540}{3000}=90.18"

"break-even quantity=\frac{FC}{P-VCunit}={432 000}{900-90.18}=533.45"

1.2

"marginal income ratio=\frac{809.82}{900}=0.8998"

"TS=\frac{FC+TP}{Cm Ratio}=\frac{432 000+600 000}{0.8998}=1 146 921.54"

1.3

after change:

"Profit=3300\times860-(282000+174 000+540+283 800)=2 838 000-740340=2 097660"

more profit after more changes, therefore we accept the offer of the manager


before change:

"Profit=3000\times900-(282000+150 000+540 000+270 000)=2 700 000-702 540=1 997 460"